OpenStudy (linn99123):

NEED HELP FAST WILL GIVE MEDAL AND FAN Twenty stamps cost $8.40. How much does it cost to buy 5 stamps? A. $42.00 B. $11.90 C. $4.20 D. $2.10

OpenStudy (anonymous):

8.40/20=?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Don't give direct answers.

OpenStudy (anonymous):

You should help the person learn.

OpenStudy (anonymous):

8.40/20= .42 .42*5=2.10 So the answer would be D, do you get how I got there.

OpenStudy (linn99123):

yea thank u for breaking it down

OpenStudy (anonymous):

Your Welcome.